A population consists of the five numbers: 1, 3, 5, 7, and 9.


Consider samples of size 3 that can be drawn from this population



2 points for every Sample and 1 point for the corresponding Mean of that Sample

1
Expert's answer
2021-11-03T18:50:02-0400

1. Mean



"\mu=\dfrac{1+3+5+7+9}{5}=5"

Variance



"\sigma^2=\dfrac{1}{5}\big((2-3.4)^2+(6-2.4)^2+(8-3.4)^2""(0-3.4)^2+(1-3.4)^2\big)=9.44"


Standard deviation



"\sigma=\sqrt{\sigma^2}=\sqrt{9.44}\approx3.0725"



2. We have population values "2,6,8,0,1" population size "N=5" and sample size "n=2." Thus, the number of possible samples which can be drawn without replacement is



"\dbinom{N}{n}=\dbinom{5}{2}=10""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,6 & 4 \\
 \hdashline
 2 & 2,8 & 5 \\
 \hdashline
 3 & 2,0 & 1 \\
 \hdashline
 4 & 2,1 & 1.5 \\
 \hdashline
 5 & 6,8 & 7 \\
 \hdashline
 6 & 6,0 & 3 \\
 \hdashline
 7 & 6,1 & 3.5 \\
 \hdashline
 8 & 8,0 & 4 \\
 \hdashline
 9 & 8,1 & 4.5 \\
 \hdashline
 10 & 0,1 & 0.5 \\ \hline
\end{array}"

3. The sampling distribution of the sample means.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 1/2 & 1& 1/10 & 1/20 & 1/40 \\
 \hdashline
 1 & 1 & 1/10 & 2/20 & 4/40 \\
 \hdashline
 3/2 & 1 & 1/10 & 3/20 & 9/40 \\
 \hdashline
 3 & 1 & 1/10 & 6/20 & 36/40 \\
 \hdashline
 7/2 & 1 & 1/10 & 7/20 & 49/40 \\
 \hdashline
 4 & 2 & 2/10 & 16/20 & 128/40 \\
 \hdashline
 9/2 & 1& 1/10 & 9/20 & 81/40 \\
 \hdashline
 5 & 1& 1/10 & 10/20 & 100/40 \\
 \hdashline
 7 & 1 & 1/10 & 14/20 & 196/40 \\
 \hdashline
 Total & 10 & 1 & 68/20 & 604/40 \\ \hline
\end{array}"

4.



"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{68}{20}=3.4"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=3.4=\mu"

5.


"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2""=\dfrac{604}{40}-(\dfrac{68}{20})^2=\dfrac{1416}{400}=\dfrac{354}{100}=3.54""\sqrt{Var(\bar{X})}=\sqrt{\dfrac{354}{100}}\approx1.8815"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{9.44}{2}(\dfrac{5-2}{5-1})""=3.54, True"